A balanced three phase load has per phase impendance of 25+j40 ohms if 400v three phase supply is connected to the load determine the

Phase current

Line current

Power factor

Power supplied to the load


1
Expert's answer
2021-02-11T08:56:01-0500

Phase impedance: "Z_P = \large\frac{Phase \space Voltage}{Phase \space Current}"

We have "Z_p = 25+40j \space [omh]" "U = 400V"

"I_p = \large\frac{U}{Z_p\sqrt{3}}" "= \large\frac{400}{40\sqrt{3}}" "= 5.78A" (Phase Current)

"I_L = \sqrt{3}I_P = 10A" (Line Current)

"cos \varphi= \large\frac{Rp}{Zp}" "= \large\frac{25}{40}" = 0.8125 (Power factor)

"P = \sqrt{3}U_LI_Lcos\varphi = \sqrt{3}*10*400*0.8125 = 2500\sqrt{3}" W